# Diverse Part Synthesis for 3D Shape Creation

### v1

@@ 3.1 Part composition network (PCN)
The PCN takes a voxelized shape part as input and outputs the predicted affine transformation parameters of the part, i.e., the scale and translation, and the transformed shape part. The PCN consists of two training units: a volumetric AE that learns a latent representation for each shape part and a spatial transformer network (STN) jaderberg2015spatial that first predicts the affine transformation parameters of parts and then scales and translates the reconstructed parts to their target positions to compose a coherent shape. Note that we do not include rotations in the transformation since the training data is aligned consistently.
@@
The training objective of the PCN can be decomposed into minimizing two losses, similar to the scheme adopted by Hu et al. hu2018predictive : (i) the reconstruction loss of the AE, i.e., the mean squared error (MSE) between the input and reconstructed shape parts $\mathbf{P}_{i}$ and $\hat{\mathbf{P}}_{i}$, and (ii) the STN loss, i.e., the MSE between the ground truth (GT) and predicted transformed shape parts ${\mathbf{P}^{\prime}}_{i}$ and $\hat{\mathbf{P}^{\prime}}_{i}$, added to the difference between the GT transformation parameters ($s_{i}$ and $\mathbf{t}_{i}$) and learned parameters ($\hat{s}_{i}$ and $\hat{\mathbf{t}}_{i}$). Therefore, the loss function of the PCN is a linear combination of the aforementioned losses, which for each training sample is written as:
@@ 3.2 Part suggestion network (PSN)
Given a partial assembly comprising one or more shape parts, the PSN predicts multiple latent vectors, where each vector corresponds to a suggested shape part that can be connected to the current partial assembly. Our PSN consists of two components: a pretrained encoder network that encodes the partial assemblies into latent codes and a deep generative model that enables multimodal part suggestions. We experiment with various deep generative models with multimodality in the implementation of the PSN, including the MDN, conditional GAN (cGAN), conditional IMLE (cIMLE), and conditional DDPM (cDDPM).
@@ 3.2.1 Mixture density network (MDN)
where $h$ is the number of distributions in the mixture, $w_{j}$ represents the mixing coefficient of the $j$-th kernel, and $\phi_{j}(\mathbf{z}_{i}\mid\mathbf{y}_{i})$ is the probability density function (PDF) of the target $\mathbf{z}_{i}$ for the $j$-th kernel. Thus, the loss function of the network is defined as minimizing the negative logarithm of the PDFs of the kernels, i.e.,
@@ 3.2.3 Conditional IMLE (cIMLE)
$\hat{\theta}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\{\hat{\mathbf{x}}_{j}\}^{m}_{j=1}}\left[\sum_{i=1}^{n}\min_{j=1}^{m}d(\mathbf{x}_{i},\hat{\mathbf{x}}_{j})\right],$ (6)
@@
The network is trained to minimize the MSE between the GT latent vector $\mathbf{z}_{i}$ and its nearest generated sample $\hat{\mathbf{z}}_{i,k}$ according to the $k$-th noise $\mathbf{n}_{i,k}$:
@@ 3.2.4 Conditional DDPM (cDDPM)
$q(\mathbf{x}^{1},\dots,\mathbf{x}^{T}\mid\mathbf{x}^{0})=\prod_{t=1}^{T}q(\mathbf{x}^{t}\mid\mathbf{x}^{t-1}),$ (10)
@@
Given a part latent code $\mathbf{z}_{i}$ and its corresponding noise $\mathbf{z}_{i}^{t}$ obtained after applying the diffusion process $t$ times, we use a U-Net ronneberger2015unet as the backbone model to learn the mapping from $\mathbf{z}_{i}^{t}$ to $\mathbf{z}_{i}^{t-1}$ at each time step $t$. The U-Net consists of a contraction path as an encoder and an expansion path as a decoder. In our U-Net structure, the contraction path has $4$ blocks of convolutional layers that performs downsampling. The expansion path, symmetrically, uses $4$ layer blocks performing upsampling. The U-Net is time-conditional by taking an additional input $t$. In our implementation, the U-Net is also conditioned on the partial assembly code $\mathbf{y}_{i}$, by concatenating $\mathbf{y}_{i}$ to $\mathbf{z}_{i}^{t}$.
@@
We denote the U-Net model as $\epsilon_{\theta}(\cdot,\cdot,\cdot)$, based on the reparameterization ho2020denoising :
@@
$\bar{\alpha}_{t}=\prod_{s=1}^{t}(1-\beta_{s}).$ (14)
We adopt the simplified training objective of DDPMs as introduced by Ho et al. ho2020denoising in our cDDPM:
$\mathcal{L}_{\text{cDDPM}}=\operatorname{MSE}(\epsilon,\epsilon_{\theta}(t,\mathbf{y}_{i},\mathbf{z}_{i}^{t})),$ (15)
@@ 4 Experimental setup
In this section, we introduce our experimental methodology.
@@
In our experiments, we use $500$ chairs and $500$ airplanes from the ShapeNet dataset with semantic segmentation chang2015shapeNet ; yi2016scalable . We also experiment with training our networks with synthetic shapes generated through a procedural modeling library (PML) ali2021evaluation . We create a set of $500$ chairs using the PML. All the categories of shapes in our experiments are composed of $3$ or $4$ semantic parts. To better evaluate the quality of the synthesized shapes, we partition each category into training and test sets according to a $4:1$ split.
@@
All the neural networks employed in our experiments are trained using the Adam optimizer kingma2015adam . For the PCN, we set the learning rate of the AE to $10^{-4}$ and the learning rate of the STN to $10^{-6}$. We employ a two-stage training scenario for the PCN, where the AE and STN are first updated simultaneously for $1000$ epochs, and then we train the STN for another $500$ epochs. We employ different training settings for different versions of PSNs. We train the MDN for $1000$ epochs using a learning rate of $10^{-4}$, the cGAN for $2000$ epochs using a learning rate of $10^{-5}$, the cIMLE for $500$ epochs using a learning rate of $10^{-4}$, and the cDDPM for $5\times 10^{5}$ epochs using a learning rate of $8\times 10^{-5}$.
@@ 5.1 Comparison of multimodal generative models
We first present a comparative study of the four different implementations of the PSN.
(a) MDN
(b) cGAN
(c) cIMLE
(d) cDDPM
Figure 4 gives representative examples of the parts suggested by the different methods for the same initial part. The initial parts are, from top to bottom, chair arms from the PML chairs, chair legs from ShapeNet chairs, and airplane wings from ShapeNet airplanes. We observe that the parts sampled from the MDN are blurry and noisy, and no diversity is observed in the cGAN-generated parts (possibly due to mode collapse). However, both cIMLE and cDDPM yield suggested parts with good visual quality and variety, with cIMLE providing slightly better semantic diversity in the generated parts. In the examples of chair arms and airplane wings, the part suggestions generated by cIMLE exhibit variations across all possible semantic labels of chair and airplane parts, while cDDPM is only able to generate chair seats and backs and airplane bodies and engines, respectively.
@@
MDN $29.07$ $37.39$ $25.67$
cGAN $10.02$ $8.966$ $6.430$
cIMLE $\mathbf{44.02}$ $\mathbf{53.84}$ $\mathbf{32.02}$
cDDPM $39.46$ $46.38$ $24.17$
@@
Since cIMLE and cDDPM perform considerably better than the other two methods, we further compare both the diversity and quality of the suggested parts generated by cIMLE and cDDPM with a more detailed quantitative evaluation involving visual inspection of the results. We first choose $100$ initial parts from each category and use cIMLE and cDDPM to generate $4$ suggested parts for each initial part. Then, we visually inspect the generated parts and determine the following measures: visual diversity (VD), semantic diversity (SD), visual quality (VQ), and positioning quality (PQ). VD denotes the count of visually diverse part instances, SD stands for the number of semantic labels associated with the parts, VQ represents the number of parts that have high visual quality, and PQ indicates the count of parts that are positioned reasonably well in relation to the initial part. Since $4$ parts were generated, the counts can lie between 00 and $4$. We present in Table 2 the averages of the aforementioned metrics for the suggested parts generated by each method across all datasets. We observe that cIMLE provides suggested parts with better diversity while the parts generated by cDDPM possess higher quality.
5.2 Shape synthesis
@@
Incremental part suggestion
Figure 5 and Figure 6 showcase example shapes produced by our method in an incremental manner based on different training data, employing the best-performing methods cIMLE and cDDPM as implementations of the PSN, respectively. We illustrate the creation of the synthesized shapes as a spanning tree, where the root node represents the initial part from which the synthesis process starts, each inner node simulates the user’s selection of a part suggestion, and the leaf nodes represent the shapes synthesized on different spanning paths. Specifically, to synthesize these shapes, we first randomly sample a latent vector from the latent space learned by the PCN and then decode the latent vector into the initial part using the implicit decoder; see the initial parts in gray in Figure 5 and Figure 6. In an interactive setting, the user can handpick the initial part. Then, our system starts to provide part suggestions in an iterative manner. Given the current shape part(s), the PSN generates multiple latent samples that are reconstructed into suggested parts by the implicit decoder. The user manually selects suggested parts that are reasonable, and our system connects the selected part to the current part(s). The newly assembled parts are given to our system for generating part suggestions in the next iteration. In our experiments, the iterative synthesis process ends after a predefined number of iterations, although in an interactive setting the user can end the process earlier when satisfied with the synthesized shape.
@@
Our iterative part suggestion also allows synthesizing shapes with edited initial parts. Figure 7 shows shapes synthesized starting from edited versions of the chair seats given in Figure 4(a), where we multiply the width of the initial part by $0.5$ in Figure 6(a) and by $0.75$ in Figure 6(b). We see that our method can generate parts that adapt to the user’s modification of the initial parts, e.g., the suggested chair backs and arms have similar widths as the initial parts and are placed at reasonable positions. Note that we base this experiment on the synthetic shapes generated by the PML, as the synthetic training data with gradual changes in shape attributes enables the learning of a smoother and denser latent manifold ali2021evaluation which allows the PSN to generate more meaningful part suggestions.
@@
We qualitatively and quantitatively compare our synthesized results to the shapes generated by two state-of-the-art 3D generative models, namely, IM-NET chen2019learning and PQ-NET wu2020pq . We compare to these two methods as they are representatives of: (i) a method that synthesizes shapes as a whole and (ii) a method that synthesizes shapes in a part-based representation. We use their released code and train their models using our data. All the results are sampled under the resolution of $64\times 64\times 64$. Both IM-NET and PQ-NET perform shape generation by randomly sampling from the latent space learned by a latent GAN. The latent GAN of IM-NET is trained with the latent vectors encoded from the entire shapes, while the latent space of PQ-NET is based on a part AE. More similarly to our method, PQ-NET further performs part generation and assembly based on recurrent neural networks.
@@
Figure 8 visually compares our results to those of IM-NET and PQ-NET, where we manually selected reasonable samples for each method. Our user-in-the-loop shape synthesis method possesses several advantages over the GAN-based generative methods. For example, the user can rule out unreasonable shape parts during their selection, and thus the shapes synthesized by our method have more visually plausible parts, while it is difficult to ensure a perfect shape with the GANs given that shapes are generated as a whole. For example, randomly sampling with the GANs can produce results with disconnected chair legs (Figure 7(a)) or floating blobs on the side of the airplane bodies or around chair legs (Figure 7(b)). Moreover, with the better reconstruction quality, our synthesized shapes preserve more geometric details in the generated parts.
@@
PML chairs IM-NET $52\,\%$ $9.241$ $5.299$
PQ-NET $19\,\%$ $46.49$ $24.97$
Ours (cIMLE) $\mathbf{55\,\%}$ $\mathbf{7.654}$ $\mathbf{2.869}$
Ours (cDDPM) $53\,\%$ $8.032$ $3.081$
ShapeNet chairs IM-NET $41\,\%$ $11.76$ $5.917$
PQ-NET $20\,\%$ $60.66$ $29.80$
Ours (cIMLE) $47\,\%$ $\mathbf{10.19}$ $3.898$
Ours (cDDPM) $\mathbf{48\,\%}$ $11.08$ $\mathbf{3.699}$
ShapeNet airplanes IM-NET $48\,\%$ $4.909$ $6.147$
PQ-NET $20\,\%$ $89.21$ $51.91$
Ours (cIMLE) $52\,\%$ $5.054$ $6.098$
Ours (cDDPM) $\mathbf{54\,\%}$ $\mathbf{4.713}$ $\mathbf{5.722}$
To quantitatively evaluate the results, we adopt the coverage (COV), minimum matching distance (MMD), and Jensen-Shannon divergence (JSD) as evaluation metrics achlioptas2018learning , where COV and JSD evaluate the diversity of the results, and MMD evaluates the fidelity. In their calculation, we use CD as the difference metric. For each method, we collect a set of shapes having $100$ samples and compare to the test set using the aforementioned metrics. For each shape, we reconstruct a corresponding triangle mesh and sample $2048$ points on the mesh surface to compute CD. Table 3 details this quantitative evaluation. Our method consistently outperforms others across all metrics and shape categories. The performance of our method varies when employing different models for implementing the PSN. When using cIMLE as the implementation of the PSN, our method achieves the highest diversity and fidelity in synthesizing shapes from the PML chairs and the highest fidelity on the ShapeNet chairs. When using cDDPM, our method attains the highest diversity and fidelity on the ShapeNet airplanes and the highest diversity on the ShapeNet chairs. We note that PQ-NET, which also operates at the part level for shape synthesis, demonstrates very poor performance in all metrics. We hypothesize that the poor results arise from our relatively small training set, lowering PQ-NET’s ability to generate high-quality parts and potentially leading to the repetition of certain parts in different shapes.
@@
Comparison to ANISE petrov2023anise
Additionally, we conduct a qualitative comparison between our method and a recent method proposed by Petrov et al. petrov2023anise , named ANISE, that focuses on reconstructing shapes as implicit fields in a part-based fashion, while also allowing a limited form of shape synthesis through shape completion and part editing. We present this comparison in Figure 9, where ANISE enables shape completion by producing a particular part to complement the given partial assembly, reconstructing a specific query shape (the first column in Figure 8(a)). It then allows users to interpolate the part to edit the reconstruction (the second column in Figure 8(a)). On the other hand, our method is able to directly provide diverse part suggestions according to the partial shape, without the need of a reference shape (Figure 8(b)). In this comparison, the implementation of the PSN is based on cDDPM, which enables the sampling of a variable number of part latent codes and thus maximizes the diversity of the generated parts.

### v2

@@ 3.1 Part composition network (PCN)
The PCN takes a voxelized shape part as input and outputs the predicted affine transformation parameters of the part, i.e., the scale and translation, and the transformed shape part. The PCN consists of two training units: a volumetric AE that learns a latent representation for each shape part and a spatial transformer network (STN) [22] that first predicts the affine transformation parameters of parts and then scales and translates the reconstructed parts to their target positions to compose a coherent shape. Note that we do not include rotations in the transformation since the training data is aligned consistently.
@@
The training objective of the PCN can be decomposed into minimizing two losses, similar to the scheme adopted by Hu et al. [55]: (i) the reconstruction loss of the AE, i.e., the mean squared error (MSE) between the input and reconstructed shape parts $\mathbf{P}_{i}$ and $\hat{\mathbf{P}}_{i}$, and (ii) the STN loss, i.e., the MSE between the ground truth (GT) and predicted transformed shape parts $\mathbf{P}^{\prime}_{i}$ and $\hat{\mathbf{P}}^{\prime}_{i}$, added to the difference between the GT transformation parameters ($s_{i}$ and $\mathbf{t}_{i}$) and learned parameters ($\hat{s}_{i}$ and $\hat{\mathbf{t}}_{i}$). Therefore, the loss function of the PCN is a linear combination of the aforementioned losses, which for each training sample is written as:
@@ 3.2 Part suggestion network (PSN)
Given a partial assembly comprising one or more shape parts, the PSN predicts multiple latent vectors, where each vector corresponds to a suggested shape part that can be connected to the current partial assembly. Our PSN consists of two components: the encoder network pretrained in the PCN that encodes the partial assemblies into latent codes and a deep generative model that enables multimodal part suggestions. We experiment with various deep generative models with multimodality in the implementation of the PSN, including the MDN, conditional GAN (cGAN), conditional IMLE (cIMLE), and conditional DDPM (cDDPM).
@@ 3.2.1 Mixture density network (MDN)
where $h$ is the number of distributions in the mixture, $w_{j}$ represents the mixing coefficient of the $j$th kernel, and $\phi_{j}(\mathbf{z}_{i}\mid\mathbf{y}_{i})$ is the probability density function (PDF) of the target $\mathbf{z}_{i}$ for the $j$th kernel. Thus, the loss function of the network is defined as minimizing the negative logarithm of the PDFs of the kernels, i.e.,
@@ 3.2.3 Conditional IMLE (cIMLE)
$\hat{\theta}=\operatorname*{arg\,min}_{\theta}\mathbb{E}_{\{\hat{\mathbf{x}}_{j}\}^{m}_{j=1}}\left[\sum^{n}_{i=1}\min^{m}_{j=1}d(\mathbf{x}_{i},\hat{\mathbf{x}}_{j})\right],$ (6)
@@
The network is trained to minimize the MSE between the GT latent vector $\mathbf{z}_{i}$ and its nearest generated sample $\hat{\mathbf{z}}_{i,k}$ according to the $k$th noise $\mathbf{n}_{i,k}$:
@@ 3.2.4 Conditional DDPM (cDDPM)
$q(\mathbf{x}^{1},\dots,\mathbf{x}^{T}\mid\mathbf{x}^{0})=\prod^{T}_{t=1}q(\mathbf{x}^{t}\mid\mathbf{x}^{t-1}),$ (10)
@@
Given a part latent code $\mathbf{z}_{i}$ and its corresponding noise $\mathbf{z}^{t}_{i}$ obtained after applying the diffusion process $t$ times, we use a U-Net [56] as the backbone model to learn the mapping from $\mathbf{z}^{t}_{i}$ to $\mathbf{z}^{t-1}_{i}$ at each time step $t$. The U-Net consists of a contraction path as an encoder and an expansion path as a decoder. In our U-Net structure, the contraction path has $4$ blocks of convolutional layers that performs downsampling. The expansion path, symmetrically, uses $4$ layer blocks performing upsampling. The U-Net is time-conditional by taking an additional input $t$. In our implementation, the U-Net is also conditioned on the partial assembly code $\mathbf{y}_{i}$, by concatenating $\mathbf{y}_{i}$ to $\mathbf{z}^{t}_{i}$.
@@
We denote the U-Net model as $\epsilon_{\theta}(\cdot,\cdot,\cdot)$, based on the reparameterization [21]:
@@
$\bar{\alpha}_{t}=\prod^{t}_{r=1}(1-\beta_{r}).$ (14)
We adopt the simplified training objective of DDPMs as introduced by Ho et al. [21] in our cDDPM:
$\mathcal{L}_{\text{cDDPM}}=\operatorname{MSE}(\epsilon,\epsilon_{\theta}(t,\mathbf{y}_{i},\mathbf{z}^{t}_{i})),$ (15)
@@ 4 Experimental setup
In this section, we introduce our experimental methodology. Implementation details of our experiments can be found in our code [^1] https://github.com/IsaacGuan/Diverse-Part-Synthesis.
@@
In our experiments, we use $500$ chairs, $500$ airplanes, and $500$ guitars from the ShapeNet dataset with semantic segmentation [57, 58]. We also experiment with training our networks with synthetic shapes generated through a procedural modeling library (PML) [59]. We create a set of $500$ chairs using the PML. All the chairs and airplanes in our experiments are composed of $3$ or $4$ semantic parts, and all the guitars are composed of $3$ semantic parts. To better evaluate the quality of the synthesized shapes, we partition each category into training and test sets according to a $4:1$ split. Note that, in general, our method is suitable for shapes that can be decomposed into semantic parts, such as man-made shapes.
@@
All the neural networks employed in our experiments are trained using the Adam optimizer [60]. For the PCN, we experimentally set the learning rate of the AE to $10^{-4}$ and the learning rate of the STN to $10^{-6}$. We employ a two-stage training scenario for the PCN, where the AE and STN are first updated simultaneously for $1000$ epochs, and then we train the STN for another $500$ epochs. We employ different training settings for different versions of PSNs. We train the MDN for $1000$ epochs using a learning rate of $10^{-4}$, the cGAN for $2000$ epochs using a learning rate of $10^{-5}$, the cIMLE for $500$ epochs using a learning rate of $10^{-4}$, and the cDDPM for $5\times 10^{5}$ epochs using a learning rate of $8\times 10^{-5}$.
@@ 5.1 Comparison of multimodal generative models
We first present a comparative study of the different implementations of the PSN.
Figure 4 gives representative examples of the parts suggested by the different methods for the same initial part. The initial parts are, from top to bottom, chair arms from the PML chairs, chair legs from the ShapeNet chairs, airplane wings from the ShapeNet airplanes, and a guitar neck from the ShapeNet guitars. We observe that the parts sampled from the MDN are blurry and noisy, and no diversity is observed in the cGAN-generated parts (possibly due to mode collapse). However, both cIMLE and cDDPM yield suggested parts with good visual quality and variety, with cIMLE providing slightly better semantic diversity in the generated parts. In the examples of chair arms and airplane wings, the part suggestions generated by cIMLE exhibit variations across all possible semantic labels of chair and airplane parts, while cDDPM is only able to generate chair seats and backs and airplane bodies and engines, respectively.
@@
MDN $27.19$ $33.87$ $22.01$
cGAN $7.546$ $6.762$ $4.823$
cIMLE $\mathbf{40.20}$ $\mathbf{47.33}$ $\mathbf{27.83}$
cDDPM $38.12$ $40.90$ $22.53$
@@
ShapeNet guitars cIMLE $\mathbf{3.42}$ $\mathbf{1.92}$ $2.77$ $2.95$
cDDPM $2.21$ $1.49$ $\mathbf{3.18}$ $\mathbf{3.65}$
@@
Since cIMLE and cDDPM perform considerably better than the other two methods, we further compare both the diversity and quality of the suggested parts generated by cIMLE and cDDPM with a more detailed quantitative evaluation involving visual inspection of the results. We first choose $100$ initial parts from each category and use cIMLE and cDDPM to generate $4$ suggested parts for each initial part. Then, we visually inspect the generated parts and determine the following measures: visual diversity (VD), semantic diversity (SD), visual quality (VQ), and positioning quality (PQ). VD denotes the count of visually diverse part instances, SD stands for the number of semantic labels associated with the parts, VQ represents the number of parts that have high visual quality, and PQ indicates the count of parts that are positioned reasonably well in relation to the initial part. Since $4$ parts were generated, the counts can lie between 00 and $4$. We present in Table 2 the averages of the aforementioned metrics for the suggested parts generated by each method across all datasets. We observe that, among the $4$ suggested parts, cIMLE provides parts with better diversity, while the parts generated by cDDPM possess higher quality.
5.2 Shape synthesis
@@
Incremental part suggestion
Figure 5 and Figure 6 showcase example shapes produced by our method in an incremental manner based on different training data (the PML chairs, ShapeNet chairs, and ShapeNet airplanes), employing the best-performing methods cIMLE and cDDPM as implementations of the PSN, respectively. We illustrate the creation of the synthesized shapes as a spanning tree, where the root node represents the initial part from which the synthesis process starts, each inner node simulates the user’s selection of a part suggestion, and the leaf nodes represent the shapes synthesized on different spanning paths. Specifically, to synthesize these shapes, we first randomly sample a latent vector from the latent space learned by the PCN and then decode the latent vector into the initial part using the implicit decoder; see the initial parts in gray in Figure 5 and Figure 6. In an interactive setting, the user can handpick the initial part. Then, our system starts to provide part suggestions in an iterative manner. Given the current shape part(s), the PSN generates multiple latent samples that are reconstructed into suggested parts by the implicit decoder. The user manually selects suggested parts that are reasonable, and our system connects the selected part to the current part(s). The newly assembled parts are given to our system for generating part suggestions in the next iteration. In our experiments, the iterative synthesis process ends after a predefined number of iterations, although in an interactive setting the user can end the process earlier when satisfied with the synthesized shape.
Our iterative part suggestion also allows synthesizing shapes with edited initial parts. Figure 7 shows shapes synthesized starting from edited versions of the chair seats given in Figure 5(a), where we multiply the width of the initial part by $0.5$ in Figure 7(a) and by $0.75$ in Figure 7(b). We see that our method can generate parts that adapt to the user’s modification of the initial parts, e.g., the suggested chair backs and arms have similar widths as the initial parts and are placed at reasonable positions. Note that we base this experiment on the synthetic shapes generated by the PML, as the synthetic training data with gradual changes in shape attributes enables the learning of a smoother and denser latent manifold [59] which allows the PSN to generate more meaningful part suggestions.
@@
We qualitatively and quantitatively compare our synthesized results to the shapes generated by two state-of-the-art 3D generative models, namely, IM-NET [1] and PQ-NET [13]. We compare to these two methods as they are representatives of: (i) a method that synthesizes shapes as a whole and (ii) a method that synthesizes shapes in a part-based representation. We use their released code and train their models using our data. All the results are sampled under the resolution of $64\times 64\times 64$. Both IM-NET and PQ-NET perform shape generation by randomly sampling from the latent space learned by a latent GAN. The latent GAN of IM-NET is trained with the latent vectors encoded from the entire shapes, while the latent space of PQ-NET is based on a part AE. More similarly to our method, PQ-NET further performs part generation and assembly based on recurrent neural networks.
@@
Figure 8 visually compares our results to those of IM-NET and PQ-NET, where we manually selected reasonable samples for each method. In each subfigure, we present from top to bottom shapes synthesized based on the PML chairs, ShapeNet chairs, ShapeNet airplanes, and ShapeNet guitars. Our user-in-the-loop shape synthesis method possesses several advantages over the GAN-based generative methods. For example, the user can rule out unreasonable shape parts during their selection, and thus the shapes synthesized by our method have more visually plausible parts, while it is difficult to ensure a perfect shape with the GANs given that shapes are generated as a whole. For example, randomly sampling with the GANs can produce results with disconnected chair legs (Figure 8(a)) or floating blobs on the side of the airplane bodies or around chair legs (Figure 8(b)). Moreover, with the better reconstruction quality, our synthesized shapes preserve more geometric details in the generated parts.
To quantitatively evaluate the results, we adopt the coverage (COV), minimum matching distance (MMD), and Jensen–Shannon divergence (JSD) as evaluation metrics [62], where COV and JSD evaluate the diversity of the results, and MMD evaluates the fidelity. In the calculation of COV and MMD, we use CD as the difference metric, and to enhance the distinction of MMD values, we base the computation of CD on the squared ED computed between each pair of points. For each method, we collect a set of shapes having $100$ samples and compare to the test set using the aforementioned metrics. For each shape, we reconstruct a corresponding triangle mesh and sample $2048$ points on the mesh surface to compute CD. Table 3 details this quantitative evaluation. Our method consistently outperforms others across all metrics and shape categories. The performance of our method varies when employing different models for implementing the PSN. When using cIMLE as the implementation of the PSN, our method achieves the highest diversity and fidelity in synthesizing shapes from the PML chairs and ShapeNet guitars, as well as the highest fidelity on the ShapeNet chairs. When using cDDPM, our method attains the highest diversity and fidelity on the ShapeNet airplanes and the highest diversity on the ShapeNet chairs. We note that PQ-NET, which also operates at the part level for shape synthesis, demonstrates very poor performance in all metrics. We hypothesize that the poor results arise from our relatively small training set, lowering PQ-NET’s ability to generate high-quality parts and potentially leading to the repetition of certain parts in different shapes.
@@
[…]
@@
Comparison to ANISE [16]
Additionally, we conduct a qualitative comparison between our method and a recent method proposed by Petrov et al. [16], named ANISE, that focuses on reconstructing shapes as implicit fields in a part-based fashion, while also allowing a limited form of shape synthesis through shape completion and part editing. We present this comparison in Figure 9, where ANISE enables shape completion by producing a particular part to complement the given partial assembly, reconstructing a specific query shape (the first column in Figure 9(a)). It then allows users to interpolate the part to edit the reconstruction (the second column in Figure 9(a)). On the other hand, our method is able to directly provide diverse part suggestions according to the partial shape, without the need of a reference shape (Figure 9(b)). In this comparison, the implementation of the PSN is based on cDDPM, which enables the sampling of a variable number of part latent codes and thus maximizes the diversity of the generated parts.